OpenStudy (anonymous):

Simplify the square root of 48s^3

OpenStudy (anonymous):

\[\sqrt{48s ^{3}}\] this?

OpenStudy (anonymous):

yeup

OpenStudy (anonymous):

48=3*2*2*2*2=3*2^4 s^3=s^2*s

OpenStudy (anonymous):

yeah...I got up to there